OpenStudy (anonymous):

Whats the largest value of x in 9x-4= sqrt(42/7x)

Parth (parthkohli):

1) Square both sides. 2) Solve the equation. 3) Choose the bigger value.

Parth (parthkohli):

\[\implies (9x - 4)^2 = {42 \over 7x} \]\[ \implies (9x - 4)^2 = {6 \over x}\]\[\implies 16 - 72 x + 81 x^2 = {6 \over x}\]\[\implies x(81x^2 - 72x + 16) = 6 \]\[\implies 81x^3 - 72x^2 + 16x = 6 \]\[\implies 81x^3 - 72x^2 + 16x - 6 = 0 \]Now solve the cubic.